OpenStudy (anonymous):

Please help me study I have to take a test in two hours, & nobody will help out. Show an equation and a solution for the problem. Arnold needs a 25% solution of nitric acid. He has 20 milliliters (ml) of a 30% solution. How many ml of a 15% solution should he add to obtain the required 25% solution?

OpenStudy (anonymous):

well if you have a test we better get busy

OpenStudy (anonymous):

you there?

OpenStudy (anonymous):

i will walk you through it if you like.

OpenStudy (anonymous):

I don't get it. & I need help.

OpenStudy (anonymous):

ok we can do it

OpenStudy (anonymous):

first off we need to find something, namely the number of ml of 15% solution to add, so we call it X so we can solve for it ok?

OpenStudy (anonymous):

you have 20 ml of a 30% solution. to this you are going to add X ml of 15% solution, so the total amount of solution will be 20 + X. the 20 you started with and the X that you added

OpenStudy (anonymous):

when you get done, you want 25% of the total to be acid, and 25% of 20 + X is written as .25(20+X)

OpenStudy (anonymous):

that will be on one side of the equal sign.

OpenStudy (anonymous):

now we have 20 ml if 30% solution and 30% of 20 is .30 * 20 = 6

OpenStudy (anonymous):

to this we are going to add X of the 15% solution and 15% of X is .15X. so the total amount of acid we will have will be 6 + .15X

OpenStudy (anonymous):

the 6 ml of acid plus the 15% of the X that we add. that will be on the other side of the equal sign. so the equation to solve is \[6+.15X=.25(20+X)\]

OpenStudy (anonymous):

how are we doing so far?

OpenStudy (anonymous):

the hard part is setting up the equation. the easier part is solving it. but before we solve it are any steps unclear?

OpenStudy (anonymous):

btw i will be happy to write the answer but since you said "i have a test" i am assuming you want to know how to do the problem

OpenStudy (anonymous):

Answer would be better.

OpenStudy (anonymous):

You asked this question twice. I answered it there.

OpenStudy (anonymous):

ok we solve \[6+.15x=.25(20+x)\] decimals are a pain so multiply everything by 100 to clear them \[600+15x=25(20+x)\] \[600+5x=500+25x\] \[600-500=25x-5x\] \[100=20x\] \[x=5\]

OpenStudy (anonymous):

notice when i cleared the decimals i had to multiply 6 by 100 as well

OpenStudy (anonymous):

@satellite73 Your answer is wrong ... your 15x turned into 5x and you ended up solving the wrong equation. Corret answer is x = 10

OpenStudy (anonymous):

@djvu Can you just write the solution & equation on her.

OpenStudy (anonymous):

Look at the image I posted.

OpenStudy (anonymous):

The equation to be solved is what satellite73 already posted : 6+.15x=.25(20+x) except he/she made an error in the calculations. Final answer is 10

OpenStudy (anonymous):

djvu you are right. somehow my 15x on one line became a 5x on the other. i should really be more careful

OpenStudy (anonymous):

Thank you.